# Applying statistical learning theory to deep learning

### v1

@@ 1.4 Deep learning in practice
In this case, our SVM model does not have a flat inductive bias, but the norm of the weights $w$ adapt to the level of complexity inherent in the data. If reality is represented by a solution with small norm, then the learning rule will achieve a solution with low complexity measure and therefore generalize. However, if we try to fit random labels, we can only fit a model with a high norm (high complexity measure), and it will fail to generalize. We can always train SVMs with zero training error $L_{S}(h)=0$. If $\exists h^{*}$ with zero generalization error $L_{\mathcal{D}}(h^{*})=0$, it will be achieved we sample complexity $|S|=O\left(||h||_{K}^{2}\right)$. Another example for this generalization is found in Minimum Description Length (MDL): A program optimized for its length $\arg\min|\text{program}|$ with $L_{S}(\text{program})=0$, is able to achieve a generalization error $L(MDL(S))\leq O\left(\frac{|\text{program}|}{|S|}\right)$. That is, a short program only requires a sample complexity proportional to its length.
These examples and the ability of deep nets to generalize implies that the size of the network is not a good measure of model complexity. This is not a new idea; it was already realized in the 1990s that kernel regression works for infinitely many features, because we rely on norm for complexity control rather than the dimensionality. It was shown by P. Bartlett in 1996 [2] that the complexity of a neural network is not controlled by the number of weights but by their magnitude. In fact, neural networks have many solutions for $w$ so that $L_{S}(w)=0$, many of which have high generalization errors. These solutions tend to have high $w$-norms. However, the solutions found in practice for neural networks using gradient descent do generalize well, and tend to have small norms, even without explicitly regularizing for low norm solutions.
Where is this implicit regularization coming from? We will try to understand this in the simplest model possible - linear regression. Consider an under-constrained least squares problem $(n>m)$:
$\min_{w\in\mathbb{R}^{n}}||Aw-b||^{2}\quad,\quad A\in\mathbb{R}^{m\times n}$
In under-constrained cases there are many choices of $w$ for which the sum of squares vanish. Imagine solving this problem with gradient descent, initialized at $w=0$. Gradient descent will definitely succeed, as this is a convex problem, and find a vanishing solution, but which solution?
@@
Gradient descent (or SGD, conjugate gradient descent, BFGS) will converge to the least norm solution $\min_{Aw=b}||w||_{2}$. The proof follows from the iterates always being spanned by the rows of $A$ (more details soon).
While we did not tell the algorithm to prefer solutions with small norms, it does in fact find the solution that minimizes this norm. This implicit regularization comes directly through the optimization process. In general we find that the optimization algorithm minimizes some norm or complexity measure, but which complexity measure?
@@ 2 Lecture 2 : Implicit bias and benign overfitting
2.1 Finishing lecture 1
To review, the high-level questions we’re trying to answer are:
@@ 2.1 Finishing lecture 1
    How much of what we’ve seen so far (in Lecture 1) fits within our classic understanding of statistical learning?
@@
One thing we’ve seen so far is that huge models (so large that they could fit even random labels) can still generalize. This is fine; it’s the same type of behavior we get with something like a hard-margin SVM or a minimum-norm predictor. What’s going on in these cases is that the $x$-axis is not actually capacity. The real measure of complexity is some kind of norm.
@@
What is this norm? Well, we are really abusing the term “norm”; what we really mean is some measure of scale that might not satisfy the rigorous definition of a norm. We can get implicit complexity control just from the optimization algorithm. For example, when we optimize an underdetermined least squares problem using gradient descent, we get the minimum-norm solution; that just comes from the optimization algorithm. So, we need to ask: what complexity measure is being minimized, and how does gradient descent minimize it?
@@
We ended last time by saying that if we change our optimization algorithm without changing the objective function, we’re actually implicitly minimizing some other complexity measure, which will change our inductive bias and thus our generalization properties. As some examples, we can compare the test performance of SGD and Path-SGD as in [25] and SGD vs. Adam as in [32]. In all cases, they reach the same final training error, but they have different final test performance values. In other words, they’re reaching different global minima of the training objective. The story we’re seeing is that the inductive bias is determined by the bias of the optimization algorithm. In other words, if we have a training loss landscape with many global minima, and we start optimizing on some hill, different optimization algorithms will move down the “hill” differently and reach different “beaches” (0 loss).
@@
An illustrative example of different optimization algorithms inducing different regularizers can be seen by studying gradient descent vs. coordinate descent. Gradient descent will get to the minimum $\ell_{2}$ norm, whereas coordinate descent will get to an approximately minimum $\ell_{1}$ norm solution. In a high-dimensional system, these two norms are extremely different; $\ell_{1}$ induces sparsity, and $\ell_{2}$ does not. This difference is significant, especially when we think about deep learning in terms of feature learning (which Nati thinks is what’s really going on). Feature selection (i.e., from a long list of features, select the relevant ones) is just sparsity, and $\ell_{1}$ regularization can achieve it. In deep learning, we want to do feature learning, not just feature selection. But what is finding new features? We can think of a continuous set of possible features, and we want to select good features from that infinite feature set. So as long as we’re not too worried about infinities, there’s not much difference between feature selection and feature learning. And you can do that better with $\ell_{1}$; $\ell_{2}$ will never give you anything that is sparse. Thus, this can be a huge difference, and all the inductive bias is coming from the algorithm here. In fact, here is the perspective on deep learning we shall take here:
@@
Deep networks can approximate any function. We kind of dismissed our universal approximation results last time because they require huge networks with seemingly unrealistic capacity. But maybe we actually are using networks that are large enough to capture all functions (since we are, at least on the data we see, able to capture all functions). So maybe we are essentially optimizing over the space of all functions. In that case, minimizing empirical error with respect to all functions doesn’t make any sense; it’s really easy to optimize the empirical error with respect to all functions, since we can just memorize the training examples and not do anything anywhere else in the domain. But in deep learning, we optimize over all functions with particular search dynamics, and although we do get to a function that has 0 training error, we don’t get to just any 0 error solution. How we optimize over the space of all functions determines which directions we like to take. Roughly speaking, we’ll get to a 0-error solution that’s “close” (to the initialization point) in some sense with respect to our geometry.
@@ 2.2 Examples
With this perspective in mind, we will now begin the main content of Lecture 2. We’ll start by examining some examples where we’re optimizing over the space of all functions, but because of the way we’re optimizing over the space of all functions, we’re actually getting good generalization.
@@ 2.2.1 Matrix completion
In some sense, it’s very easy to solve this optimization problem. We can complete the observed entries and put 0 everywhere else, but it won’t help us recover the unobserved entries. The problem is underdetermined. So what do we do? We can run gradient descent directly on $X$, or - alternatively - we can replace $X$ with $UV^{\top}$ ($U,V$ full dimensional, therefore no rank constraint on $X=UV^{\top}$) and run gradient descent on $U,V$. Figure 3 compares the results of these two procedures when the ground-truth matrix $X^{*}$ has rank 2. We also see how slight variations in gradient descent on $U,V$ change which solution we converge to. But the bigger effect is coming from the reparameterization (from $X$ to $UV^{\top}$).
So what is going on here? In this case, we have a good understanding, though not complete. The initial proposal in [12] was that gradient descent is converging to a low nuclear norm solution - i.e., nuclear norm is the relevant complexity measure that gradient descent is minimizing. We know that minimizing the nuclear norm can give you good generalization when you have an approximate low-rank matrix. Minimization of the nuclear norm is proved rigorously in some cases (e.g., under restricted isometry property (RIP) in [15]) and turns out to not always be the case (e.g., see counterexample in Example 5.9 in [16]).
@@ 2.2.2 Single overparameterized linear unit: $\mathbb{R}^{d}\to\mathbb{R}$
If we train a single unit with gradient descent using the logistic (“cross entropy”) loss, we converge to the max-margin separator (the hard-margin SVM predictor), which has an implicit $\ell_{2}$ norm in it:
@@
This holds regardless of initialization. We will go over this result in detail in a future lecture.
@@ 2.2.3 Deep linear network: $\mathbb{R}^{d}\to\mathbb{R}$
Now, we can ask what happens in a deeper network (with only linear activations). Let $w$ denote the weights of all the layers. Then our deep linear network implements the same linear mapping as above:
@@
When we run gradient descent on $w$ (vs. $\beta$, as we did above), one might think that this reparameterization could affect the search geometry. However, in this case, the inductive bias is actually the same as above:
@@ 2.2.4 Linear convolutional networks
which is still just a reparameterization of our original linear function from $\mathbb{R}^{d}\to\mathbb{R}$. Now we can ask what happens when we train this model using gradient descent.
@@ 2.2.5 Infinite-width ReLU network
This would be a very sensible penalty to choose, since it is kind of a smoothness-inducing penalty. The interesting thing is that we didn’t explicitly choose it; it came to us just from gradient descent on this parameterization.
@@
$\int|\partial_{b}^{d+1}\operatorname{Radon}(h)|.$ (9)
Once again, we get some kind of sensible smoothness penalty. This result is rigorous for logistic loss (doesn’t depend on initialization or learning rate). For squared loss, we don’t know how to analyze it exactly, although we expect something similar. See [27], [26], [7] for more details.
@@ 2.2.6 Takeaways from these examples
What have we been doing?
The game here is that we want to understand what happens in the space of functions. The inductive bias in parameter space is relatively simple ($\ell_{2}$ or something similar, often). But what we really care about is what happens in function space, which can be very rich. A large part of the optimization problem is the architecture. The classical view is that the architecture is important because it limits what functions you can get; however, that’s not the case here. The architecture is important because it determines the mapping from parameter space to function space and is the biggest contributor to the geometry by which you’re searching in function space.
@@
The answer is: sort of. You can get some asymptotic results showing that, for some restricted class of models with the logistic loss, everything boils down to $\ell_{2}$. However, we’ll also see that in many cases, this is not true, and you’ll get something very different (e.g., under squared loss, or even under logistic loss non-asymptotically).
2.3 Why the red curve goes down in Figure 3
We now understand that, in Figure 3, we have complexity control coming from the algorithm, and this is what stops us from overfitting. The main question this led us to ask is: what is this complexity control? (which we just studied, in several examples.)
But there is another thing that is going on here. What we have studied so far explains why we might be able to generalize well even at a large number of hidden units (i.e., we have complexity control coming from the algorithm, even though we’re optimizing in the space of all functions). However, we haven’t yet explained why the red curve (test error) actually goes down as the number of hidden units increases. Recall, the $x$-axis is not complexity (complexity is something else - some norm-based complexity, probably).
@@ 2.3 Why the red curve goes down in Figure 3
We can actually see similar behavior even with kernel methods. Let us consider the Gaussian kernel, which corresponds to an infinite-dimensional feature space. Let us think of what happens if we use a finite approximation to the Gaussian kernel. Concretely, we have the Gaussian kernel $\langle\phi_{\infty}(x),\phi_{\infty}(x^{\prime})\rangle=e^{-\|x-x^{\prime}\|^{2}}$ and the finite-dimensional feature mapping
$\phi_{d}(x)[i]=\frac{1}{\sqrt{d}}\cos(\langle\omega_{i},x\rangle+\theta_{i}).$
The algorithm returns $A(S)=\arg\min\|w\|$ s.t. $L_{S}\left(x\mapsto\langle w,\phi_{d}(x)\rangle\right)=0$, i.e., $\forall(x_{i},y_{i})\in S$, $y_{i}=\langle w,\phi_{d}(x_{i})\rangle$. As $d\to\infty$, we approach the Gaussian kernel. Once we have more features than data points, we can already get 0 training error. But we are not doing it with the Gaussian kernel yet; we are doing it with an approximation of the Gaussian kernel. If the RKHS norm induced by the Gaussian kernel is our “correct” complexity measure, then as $d\to\infty$, we are approximating it better and better. So we are minimizing a complexity measure that’s a better and better approximation of the complexity measure we want. So as the dimensionality increases, the test error improves.
@@
We can also see this in matrix completion. Suppose $X$ is $n\times n$, we observe $nk$ entries, and we parameterize $X$ as $UV^{\top}$, where $U,V\in\mathbb{R}^{n\times d}$ (thereby adding a rank-$d$ constraint). As above, $d$ captures the quality of our approximation. Suppose the “right” complexity measure is nuclear norm. We have two different regimes:
@@
So as we increase the rank constraint, the test error becomes better and better. As we increase the rank, we’re getting closer to what we really want, which is a full-rank low nuclear norm matrix. So $d$ is not our complexity measure; rather, it’s the dimensionality of the approximation to the infinite-dimensional system.
Commentary.
We claim that this is what’s happening in neural networks too. The real object we should be learning with is an infinite-size network. But we cannot really represent an infinite-size network? Instead, we represent a truncated, finite-dimensional representation of it. As our truncation becomes finer and finer, our representation becomes better and better. We want the size to be so big that we essentially have a good approximation to the infinite-size model. Our methods should not rely on the size of the approximation being small. So we should start by understanding the infinite size and then worry about the question: “how large do we need our model to be in order for it to be considered infinite?”.
2.4 What we don’t yet fully understand: benign overfitting
However, there’s still one thing that doesn’t really fit our understanding. And this is something we were completely blind to for many years, even though it was right in front of us. It was pointed out only fairly recently by Misha Belkin. This one thing is that we are getting good generalization even though we are insisting on a 0-training error solution in noisy situations (situations where the approximation error is nonzero). In particular, in Figure 5, we want to balance complexity and training error, so we know we want to be somewhere on this regularization frontier. But the solutions we are finding are the minimum-complexity 0-error solutions - an extreme point of the frontier, and we are seeing this even in fairly noisy cases, where we would expect to be somewhere on the frontier that strikes more of a balance between complexity and training error.

### v2

@@ 1.4 Deep learning in practice
In this case, our SVM model does not have a flat inductive bias, but the norm of the weights $w$ adapt to the level of complexity inherent in the data. If reality is represented by a solution with small norm, then the learning rule will achieve a solution with low complexity measure and therefore generalize. However, if we try to fit random labels, we can only fit a model with a high norm (high complexity measure), and it will fail to generalize. We can always train SVMs with zero training error $L_{S}(h)=0$. If $\exists h^{*}$ with zero generalization error $L_{\mathcal{D}}(h^{*})=0$, it will be achieved with sample complexity $|S|=O\left(||h||_{K}^{2}\right)$. Another example for this generalization is found in Minimum Description Length (MDL): A program optimized for its length ($\arg\min|\text{program}|$) with $L_{S}(\text{program})=0$, is able to achieve a generalization error $L(MDL(S))\leq O\left(\frac{|\text{program}|}{|S|}\right)$. That is, a short program only requires a sample complexity proportional to its length.
These examples and the ability of deep nets to generalize implies that the size of the network is not a good measure of model complexity. This is not a new idea; as it was already realized in the 1990s that kernel regression works for infinitely many features, because we rely on norm for complexity control (assuming the hypothesis class is a ball with fixed radius in the corresponding Reproducing Kernel Hilbert Space (RKHS)) rather than the dimensionality. It was shown in 1996 [4] that the complexity of a neural network is not controlled by the number of weights but by their magnitude. In fact, neural networks have many solutions for weights $w$ verifying $L_{S}(w)=0$, many of which have high generalization errors. These solutions tend to have high $w$-norms. However, the solutions found in practice for neural networks using gradient descent do generalize well, and tend to have small norms, even without explicitly regularizing for low norm solutions.
Where is this implicit regularization coming from? We will try to understand this in the simplest model possible - linear regression - in the next lecture. Consider an under-constrained least squares problem with $(n>m)$:
$\min_{w\in\mathbb{R}^{n}}||Aw-b||^{2}\quad,\quad A\in\mathbb{R}^{m\times n}.$
In under-constrained cases there are many choices of $w$ for which the sum of squares vanishes. Imagine solving this problem with gradient descent, initialized at $w=0$. Gradient descent will definitely succeed, as this is a convex problem, and find a vanishing error solution, but which solution?
@@
Gradient descent (or SGD, conjugate gradient descent, BFGS) will converge to the least $\ell_{2}$-norm solution $\min_{Aw=b}||w||_{2}$. The proof follows from the iterates always being spanned by the rows of $A$ (more details will be given in the next lecture).
While we did not explicitly design the algorithm to prefer solutions with small norm, it does in fact find the solution that minimizes this norm. This implicit regularization comes directly through the optimization process. In general we find that optimization algorithms minimize some norm or complexity measure, but which complexity measure?
@@ 2 Lecture 2 : Implicit bias and benign overfitting
2.1 Choosing the right complexity measure
The discussion iniated in the previous lecture leads to the following main questions :
@@ 2.1 Choosing the right complexity measure
    How much of what we’ve seen so far fits within our classic understanding of statistical learning?
@@
One thing we’ve seen is that huge models (so large that they could fit even random labels) can still generalize well. This does not contradict the classical understanding of supervised learning; as it’s the same type of behavior obtained with something like a hard-margin SVM or a minimum-norm predictor. The reason for this is that what actually governs the ability of the model to generalize is not actually its capacity. The real measure of complexity is some kind of norm.
@@
What is this norm? First of all, we are not using the word “norm” in an appropriate mathematical sense; what we really mean is some measure of scale that might not satisfy the rigorous definition of a norm. We can get implicit complexity control just from the optimization algorithm. For example, when we optimize an underdetermined least squares problem using gradient descent, we get the minimum-norm solution; that just comes from the optimization algorithm. So, we need to ask: what complexity measure is being minimized, and how does gradient descent minimize it?
@@
We ended the previous lecture by saying that if we change our optimization algorithm without changing the objective function, we’re actually implicitly minimizing some other complexity measure, which will change the inductive bias and thus the generalization properties of our model. As some examples, we can compare the test performance of SGD and Path-SGD as in [39] and SGD with the Adam optimizer as in [51]. In all cases, they reach the same final training error, but they have different final test performance values. In other words, they’re reaching different global minima of the training objective. Thus what we are observing is that the inductive bias is determined by the bias of the optimization algorithm. In other words, if we have a training loss landscape with many global minima, and we start optimizing on some “hill”, different optimization algorithms will move down the “hill” differently and reach different “beaches” (0 loss).
@@
An illustrative example of different optimization algorithms inducing different regularizers can be seen by studying gradient descent vs. coordinate descent. Gradient descent will get to the minimum $\ell_{2}$ norm, whereas coordinate descent will get to an approximately minimum $\ell_{1}$ norm solution. In a high-dimensional system, these two norms are extremely different; $\ell_{1}$ induces sparsity, and $\ell_{2}$ does not. This difference is significant, especially when we think about deep learning in terms of feature learning. Feature selection (i.e., from a long list of features, select the relevant ones) is just sparsity, and $\ell_{1}$ regularization can achieve it. In deep learning, we want to do feature learning, not just feature selection. But what is finding new features? We can think of a continuous set of possible features, and we want to select good features from that infinite feature set. So as long as we’re not too worried about infinities, there’s not much difference between feature selection and feature learning. This can be quantified by establishing how a givne algorithm implicitly minimized an $\ell_{1}$ norm that will induce sparsity; as opposed to an $\ell_{2}$ norm that does not. Thus, this can be a huge difference, and all the inductive bias is coming from the algorithm here. In fact, here is the perspective on deep learning we shall take here:
@@
Deep networks can approximate any function. We somewhat dismissed our universal approximation results in the previous lecture because they require huge networks with seemingly unrealistic capacity. But perhaps we actually are using networks that are large enough to capture all functions (since we are, at least on the available data, able to capture all functions). So maybe we are essentially optimizing over the space of all functions. In that case, minimizing empirical error with respect to all functions doesn’t make any sense; it’s really easy to optimize the empirical error with respect to all functions, since we can just memorize the training examples and not do anything anywhere else in the domain. But in deep learning, we optimize over all functions with particular search dynamics, and although we do get to a function that has 0 training error, we don’t get to just any 0 error solution. How we optimize over the space of all functions determines which directions we like to take. Roughly speaking, we’ll get to a 0-error solution that’s “close” to the initialization point in some sense with respect to our geometry.
@@ 2.2 Examples
With this perspective in mind, let us now discuss a few examples in which we are seemingly optimizing over the entire function class of a given problem but where the choice of optimization dynamics implicitly imposes a form of regularity on the solution, leading to good generalization.
@@ 2.2.1 Matrix completion
In some sense, it’s very easy to solve this optimization problem. We can complete the observed entries and put 0 everywhere else, but it won’t help us recover the unobserved entries. The problem is underdetermined. So what do we do? We can run gradient descent directly on $X$, or - alternatively - we can replace $X$ with $UV^{\top}$ ($U,V$ full rank, therefore no rank constraint on $X=UV^{\top}$) and run gradient descent on $U,V$. Figure 3 compares the results of these two procedures when the ground-truth matrix $X^{*}$ has rank 2. We also see how slight variations in gradient descent on $U,V$ change which solution we converge to. But the bigger effect is coming from the reparameterization (from $X$ to $UV^{\top}$).
So how can the implicit bias be understood here? In this case, we have a good understanding, though not complete. The initial proposal in [23] was that gradient descent is converging to a low nuclear norm solution - i.e., nuclear norm is the relevant complexity measure that gradient descent is minimizing. We know that minimizing the nuclear norm can give you good generalization when you have an approximate low-rank matrix. Minimization of the nuclear norm is proved rigorously in some cases (e.g., under restricted isometry property (RIP) in [29]) and turns out to not always be the case (e.g., see counterexample in Example 5.9 in [30]).
@@ 2.2.2 Single overparameterized linear unit: $\mathbb{R}^{d}\to\mathbb{R}$
If we train a single unit with gradient descent using the logistic (“cross entropy”) loss, we converge to the max-margin separator (the hard-margin SVM predictor) [47], which involves an implicit $\ell_{2}$ norm regularization :
@@
This holds regardless of initialization. We will go over this result in detail later on.
@@ 2.2.3 Deep linear network: $\mathbb{R}^{d}\to\mathbb{R}$
Now, let us consider what happens in a deeper network (with only linear activations). Let $w$ denote the weights of all the layers. Then our deep linear network implements the same linear mapping as above:
@@
When we run gradient descent on $w$ (as opposed to $\beta$, as we did above), one might think that this reparameterization could affect the search geometry. However, in this case, the inductive bias is actually the same as above:
@@ 2.2.4 Linear convolutional networks
with $10\leq d\leq D-$ 1, which is still just a reparameterization of our original linear function from $\mathbb{R}^{d}\to\mathbb{R}$. Now we can ask what happens when we train this model using gradient descent.
@@ 2.2.5 Infinite-width ReLU network
This would be a very sensible penalty to choose, since it is kind of a smoothness-inducing penalty. The interesting thing is that we didn’t explicitly choose it; it appeared solely from gradient descent on this parameterization.
@@
$\int|\partial_{b}^{d+1}\operatorname{Radon}(h)|,$ (9)
where $Radon(h)$ designates the Radon transform of $h$ with parameter $b$. Once again, we obtain a form of sensible smoothness penalty. This result is rigorous for logistic loss (doesn’t depend on initialization or learning rate). For squared loss, we don’t know how to analyze it exactly, although we expect something similar. See [44], [40], [15] for more details.
@@ 2.2.6 Takeaways from these examples
Main contributors to implicit bias
The game here is that we want to understand what happens in the space of functions. The inductive bias in parameter space is relatively simple ($\ell_{2}$ or something similar, often). But what we really care about is what happens in function space, which can be very rich. A large part of the optimization problem is dictated by the architecture. The classical view is that the architecture is important because it limits what functions may be obtained; however, that’s not the case here. The architecture is important because it determines the mapping from parameter space to function space and is the biggest contributor to the geometry by which we are searching in function space.
@@
The answer is both yes and no. It is possible to obtain some asymptotic results showing that, for some restricted class of models with the logistic loss, everything boils down to an $\ell_{2}$ norm. However, we’ll also see that in many cases, this is not true, and we’ll obtain something very different (e.g., under squared loss, or even under logistic loss non-asymptotically).
2.3 Beyond mitigating overfitting : better generaliztion through reparameterization
We now understand that, in Figure 3, we have complexity control coming from the algorithm, and this is what stops the model from overfitting. The main question this led us to ask is: what is this complexity control? (which we just studied, in several examples.)
But there is another thing that is going on here. What we have studied so far explains why we might be able to generalize well even at a large number of hidden units (i.e., we have complexity control coming from the algorithm, even though we’re optimizing in the space of all functions). However, we haven’t yet explained why the red curve in Figure 3 (test error) actually goes down as the number of hidden units increases. Recall, the $x$-axis is not complexity (complexity is something else - some norm-based complexity, probably).
@@ 2.3 Beyond mitigating overfitting : better generaliztion through reparameterization
We can actually see similar behavior even with kernel methods. Let us consider the Gaussian kernel, which corresponds to an infinite-dimensional feature space, the RKHS corresponding to this kernel, denoted $\mathcal{H}$. Let us think of what happens if we use a finite approximation to the Gaussian kernel. Concretely, we have the Gaussian kernel $\langle\phi_{\infty}(x),\phi_{\infty}(x^{\prime})\rangle=e^{-\|x-x^{\prime}\|^{2}}$ and the finite-dimensional random feature mapping
$\phi_{d}(x)[i]=\frac{1}{\sqrt{d}}\cos(\langle\omega_{i},x\rangle+\theta_{i})$
approximating $\mathcal{H}$ [42]. According to Bochner’s theorem, see e.g. [11], the Fourier transform $p(\omega)$ of any translation invariant kernel is a probability distribution. The features $\phi_{d}(x)[i]$ then correspond to the discretization of the representation of the Gaussian kernel as the inverse Fourier transform of $p(\omega)$. Thus the parameters $\omega_{i},\theta_{i}$ are not learned but drawn randomly. We refer the reader to [42] and references therein for details of implementation of this method and, to e.g. [43] for related theoretical guarantees. For a given empirical loss $L_{S}$ based on an i.i.d. sample $S$, the algorithm returns
$A(S)=\arg\min\|w\|_{\mathcal{H}}$
[…]
As $d\to\infty$, we approach the Gaussian kernel. Once we have more features than data points, we can already get 0 training error. But we are not doing it with the Gaussian kernel yet; we are doing it with an approximation of the Gaussian kernel obtained from the finite-dimensional features. If the RKHS norm induced by the Gaussian kernel is our “correct” complexity measure, then as $d\to\infty$, we are approximating it better and better. So we are minimizing a complexity measure that’s a better and better approximation of the complexity measure we want. So as the dimensionality increases, the test error improves.
@@
We can also see this in matrix completion, a finite dimensional problem. Suppose $X$ is $n\times n$, we observe $nk$ entries, and we parameterize $X$ as $UV^{\top}$, where $U,V\in\mathbb{R}^{n\times d}$ (thereby adding a rank-$d$ constraint). As above, $d$ captures the quality of our approximation. Suppose the “right” complexity measure is nuclear norm. Letting $\hat{L}$ be the empirical loss, we have two different regimes:
@@
So as we increase the rank constraint, the test error becomes better and better. As we increase the rank, we’re getting closer to what we really want, which is a full-rank low nuclear norm matrix. So $d$ is not our complexity measure; rather, it’s the dimensionality of the approximation to the full-rank system.
Remark.
We claim that this is what’s happening in neural networks too. The real object we should be learning with is an infinite-size network. But we cannot really represent an infinite-size network? Instead, we represent a truncated, finite-dimensional representation of it. As our truncation becomes finer and finer, our representation becomes better and better. We want the size to be so big that we essentially have a good approximation to the infinite-size model. Our methods should not rely on the size of the approximation being small. So we should start by understanding networks of infinite size and then worry about the question: “how large do we need our model to be in order for it to be considered infinite?”.
2.4 A more recent form of implicit bias : benign overfitting
There is still one case that doesn’t really fit the classical understanding of supervised learning, which was recently pointed out in [9, 8]. The observation is the following : we are getting good generalization even though we are insisting on a 0-training error solution in noisy situations (situations where the approximation error is nonzero). In particular, in Figure 5, we want to balance complexity and training error, so we know we want to be somewhere on this regularization frontier. But the solutions we are finding are the minimum-complexity 0-error solutions - an extreme point of the frontier, and we are seeing this even in fairly noisy cases, where we would expect to be somewhere on the frontier that strikes more of a balance between complexity and training error.